Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Ten strips of paper are numbered from 1 to 10 and placed . .

Expert replies
by Vincen » Tue Dec 19, 2017 7:06 am
Ten strips of paper are numbered from 1 to 10 and placed in a bag. If three numbers are drawn from the bag at random, what is the probability that the sum of the numbers drawn will be odd?

A. 1/12
B. 5/36
C. 15/36
D. 1/2
E. 11/18

The OA is D.

What are the equations I should set to solve this PS question? Experts, can you help me? Thanks in advanced.
Join the discussion
Source: — Problem Solving |

by [email protected] » Wed Dec 20, 2017 12:20 pm
Hi Vincen,

We're told that 10 strips of paper (numbered from 1 to 10) are placed in a bag and that three numbers are drawn from the bag at random. We're asked for the probability that the SUM of the numbers drawn will be ODD. This question can be approached as either a 'Combination' or a 'Permutation'

For the sum to be ODD, we need to draw either 1 odd and 2 evens OR 3 odds. The probabilities of those possible outcomes (written as permutations) are:

(odd)(even)(even) = (5/10)(5/9)(4/8) = 100/720
(even)(odd)(even) = (5/10)(5/9)(4/8) = 100/720
(even)(even)(odd) = (5/10)(4/9)(5/8) = 100/720
(odd)(odd)(odd) = (5/10)(4/9)(3/8) = 60/720

Total = (100/720) + (100/720) + (100/720) + (60/720) = 360/720 = 1/2

Final Answer: D

GMAT assassins aren't born, they're made,
Rich
Contact Rich at [email protected]
Image
Join the discussion

by GMATWisdom » Wed Dec 20, 2017 2:43 pm
Vincen wrote:Ten strips of paper are numbered from 1 to 10 and placed in a bag. If three numbers are drawn from the bag at random, what is the probability that the sum of the numbers drawn will be odd?

A. 1/12
B. 5/36
C. 15/36
D. 1/2
E. 11/18

The OA is D.

What are the equations I should set to solve this PS question? Experts, can you help me? Thanks in advanced.
Total number of ways of selecting groups of three numbers out of 10 is 10c3=120

There are 5 even numbers and 5 odd numbers

The sum of all the three numbers would be odd only if either 1. two numbers are even and one is odd
or if 2. all the thee numbers are odd.

when two numbers are even and one is odd the number of groups would be 5c2 x 5c1=10x5=50
when all the three numbers are odd combinations would be 5c3=10

Thus the probability of sum being odd is (50+10)/120=1/2
Hence option D
Join the discussion

by GMATGuruNY » Wed Dec 20, 2017 3:51 pm
Vincen wrote:Ten strips of paper are numbered from 1 to 10 and placed in a bag. If three numbers are drawn from the bag at random, what is the probability that the sum of the numbers drawn will be odd?

A. 1/12
B. 5/36
C. 15/36
D. 1/2
E. 11/18
Ways to get an ODD sum:
3 odd numbers
1 odd number, 2 even numbers


Ways to get an EVEN sum:
3 even numbers
1 even number, 2 odd numbers


Notice that the blue outcomes and the red outcomes are perfectly parallel.
Implication:
Since the 10 strips contain an equal number of odd numbers and even numbers, the probability of getting the blue outcomes must be equal to the probability of getting the red outcomes.
Thus:
P(odd sum) = 1/2.
P(even sum) = 1/2.

The correct answer is D.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by [email protected] » Thu Dec 21, 2017 10:14 am
Hi Vincen,

We're told that 10 strips of paper (numbered from 1 to 10) are placed in a bag and that three numbers are drawn from the bag at random. We're asked for the probability that the SUM of the numbers drawn will be ODD. This question can be approached as either a 'Combination' or a 'Permutation'

For the sum to be ODD, we need to draw either 1 odd and 2 evens OR 3 odds. The probabilities of those possible outcomes (written as permutations) are:

(odd)(even)(even) = (5/10)(5/9)(4/8) = 100/720
(even)(odd)(even) = (5/10)(5/9)(4/8) = 100/720
(even)(even)(odd) = (5/10)(4/9)(5/8) = 100/720
(odd)(odd)(odd) = (5/10)(4/9)(3/8) = 60/720

Total = (100/720) + (100/720) + (100/720) + (60/720) = 360/720 = 1/2

Final Answer: D

GMAT assassins aren't born, they're made,
Rich
Contact Rich at [email protected]
Image
Join the discussion

by Scott@TargetTestPrep » Mon Sep 09, 2019 6:29 pm
Vincen wrote:Ten strips of paper are numbered from 1 to 10 and placed in a bag. If three numbers are drawn from the bag at random, what is the probability that the sum of the numbers drawn will be odd?

A. 1/12
B. 5/36
C. 15/36
D. 1/2
E. 11/18

The OA is D.

What are the equations I should set to solve this PS question? Experts, can you help me? Thanks in advanced.

The three number drawn must be all odd OR two even and one odd.

P(3 odd) = 5C3/10C3 = [(5 x 4 x 3)/(3 x 2)]/[(10 x 9 x 8)/(3 x 2)] = (5 x 4 x 3)/(10 x 9 x 8) = 1/(2 x 3 x 2) = 1/12

P(2 even and 1 odd) = (5C2 x 5C1)/10C3 = [(5 x 4)/2 x 5]/[(10 x 9 x 8)/(3 x 2)] = [5 x 2 x 5]/[5 x 3 x 8] = 10/24 = 5/12

Therefore, the overall probability is 1/12 + 5/12 = 6/12 = 1/2.

Answer: D

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion